# Some Remarks on Positive/Negative Feedback

### v2

@@ 1 Nonlinear scalar system
For later use, we record the following.
@@
Proof.
For notational convenience, introduce the function
$F\colon\mathbb{R}^{3}\times\mathbb{R}\to\mathbb{R},~~(z,s)=\big((\rho,\xi,v),s\big)\mapsto vG(\rho,\xi,sv).$
Write $V:=[-1,-v^{*}]\cup[v^{*},1]$ and $Z:=K\times V$, a compact set. Observe that
$\chi(s)=\min_{z\in Z}F(z,s)~~\forall\,s\in\mathbb{R},$
and so, for each $s\in\mathbb{R}$, there exists $z_{s}\in Z$ such that $\chi(s)=F(z_{s},s)$. Fix $s\in\mathbb{R}$ arbitrarily. We will show that $\chi$ is continuous at $s$. Let $\varepsilon>0$. By continuity of $F$, for each $z\in Z$, there exists $\delta(z)>0$ such that, for all $\zeta,\sigma\in\mathbb{R}$ we have
$\|\zeta-z\|+|\sigma-s|<\delta(z)\implies|F(\zeta,\sigma)-F(z,s)|<\varepsilon$
and so $\cup_{z\in Z}\{(\zeta,\sigma)\mid\|\zeta-z\|+|\sigma-s|<\delta(z)\}$ is an open cover of $Z\times\{s\}$. By compactness of the latter, there exists a finite sub-cover. In particular, there exist finitely many points $z_{1},\ldots,z_{n}\in Z$ such that
$Z\times\{s\}\subset\bigcup_{i=1}^{n}\{(\zeta,\sigma)\mid\|\zeta-z_{i}\|+|\sigma-s|<\delta(z_{i})\}.$
Define $\delta^{*}:=\min\{\delta(z_{1}),\ldots,\delta(z_{n})\}$ and observe that
$\forall\,\zeta,\sigma\in\mathbb{R}\colon~~z\in Z,~|\sigma-s|<\delta^{*}\implies|F(z,\sigma)-F(z,s)|<\varepsilon.$
It follows that, for all $\sigma\in\mathbb{R}$ with $|\sigma-s|<\delta^{*}$ we have
$\varepsilon>|F(z_{s},\sigma)-F(z_{s},s)|\geq F(z_{s},\sigma)-F(z_{s},s)\geq\chi(\sigma)-\chi(s)$
and, similarly,
$\varepsilon>|F(z_{\sigma},\sigma)-F(z_{\sigma},s)|\geq F(z_{\sigma},s)-F(z_{\sigma},\sigma)\geq\chi(s)-\chi(\sigma).$
Therefore, $|\chi(\sigma)-\chi(s)|<\varepsilon$ for all $\sigma\in\mathbb{R}$ with $|\sigma-s|<\delta^{*}$, and so $\chi$ is continuous at $s$. ∎

### v3

@@ 1 Nonlinear scalar system
In essence, Definition 1.1 attempts to capture the following intuitive idea (loosely stated). The existence of a control direction is a system requirement that large input values $v$ should dominate in the sense that, with the pair $(\rho,\xi)$ of its first two arguments compactly constrained, the function $G$ is such that $|G(\rho,\xi,v)|$ can attain arbitrarily large values by taking $|v|$ sufficiently large. For example, if $G\colon(\rho,\xi,v)\mapsto f(\rho,\xi)g(v)$, then surjectivity of (continuous) $g$ is a necessary condition for the existence of a control direction.
For later use, we record the following fact (a particular consequence of a general result in variational analysis ).